consider a peak rectifierfed by 60Hz sinusoidal having a peak valueVp=100v. let the load resistance R=10Kohm calculate the fraction of cycle during which the diode is conducting


1
Expert's answer
2021-09-08T17:44:16-0400

"\theta=\sqrt{\dfrac{2V_r}{V_p}}=\sqrt{\dfrac{2\times{2}}{100}}=0.2rad\\
\dfrac{0.2}{2\pi}\times100\%=3.18\%"